# Graph Optimality-Aware Stochastic LiDAR Bundle Adjustment with Progressive Spatial Smoothing

### v1

@@ 6.3 Evaluation on Air-based Benchmark Dataset
Figure 10: The point clouds and dense graphs in MARS-LVIG (Li et al., 2024a) dataset. Edges of the graph are marked as yellow lines. The nodes of the graph are marked as green axes.
Figure 11: Absolute Position Error (APE) distribution of PSS-GOSO in MARS-LVIG (Li et al., 2024a) dataset.
6.4 Evaluation on Proprietary Large-scale Port Dataset
The 3D prior map is essential for automatic robot delivery in large-scale port environments, where the complexity of the scene and the high volume of data present significant challenges. We conduct a comprehensive evaluation using data collected from an Ouster OS1-64 laser scanner across a $400\text{m}\times 400\text{m}$ port area, supplemented by RTK-GNSS data for ground truth. Initial pose estimates are obtained using Fast-LIO. The dense graph and point clouds generated by PSS-GOSO are illustrated in Fig. 12, and the Absolute Pose Error (APE) distribution is shown in Fig. 13. The experimental results, summarized in Table 3, reveal that PSS-GOSO achieves an APE of 1.46 meters for the port UGV sequence. This performance significantly surpasses the initial APE of 8.84 meters, the APE of 5.63 meters from the pose-graph method (Koide et al., 2019), and the 3.48 meters APE from the hierarchical bundle adjustment (HBA) method (Liu et al., 2023). These results highlight PSS-GOSO’s superior accuracy and effectiveness in handling large-scale, complex environments, demonstrating its capability to improve 3D mapping and pose estimation for automated robot delivery in ports.
Table 3: Absolute Position Error (APE) of different methods in the large-scale port dataset [m]. The best results are in BOLD.
Sequence Initial (Xu et al., 2022) PSS-GOSO Pose-graph  (Koide et al., 2019) HBA (Liu et al., 2023)
Port UGV 8.84 1.46 5.63 3.48
@@ 6.4 Evaluation on Proprietary Large-scale Port Dataset
6.5 Evaluation on Proprietary Cross-platform Indoor-to-outdoor Dataset
The 3D prior map is essential for applications such as outdoor-to-indoor surveillance and elderly health care, where accurate spatial information is crucial. Large-scale mapping tasks often involve cross-platform and multi-session data fusion, posing significant challenges for existing LBA methods. Traditional sequential splitting strategies, such as those used in HBA (Liu et al., 2023), are insufficient for addressing the complexities of multi-session and cross-platform data integration. In contrast, PSS-GOSO effectively tackles these challenges by partitioning the graph based on spatial relationships, allowing it to manage cross-platform and multi-session fusion tasks more efficiently. For our evaluation, initial pose estimates are derived using Fast-LIO and manual registration of submaps from different sessions. The dense graph and point clouds generated by PSS-GOSO are shown in Fig. 14. To assess the accuracy of the point clouds produced by various methods, ground control points (GCPs) are collected using Leica RTC360 survey scanners, with their distribution displayed in Fig. 15. The average errors of the GCPs, detailed in Table 4, indicate that PSS-GOSO reduced initial errors from 3.02 meters to 0.98 meters, demonstrating superior accuracy compared to existing pose-graph-based methods. This highlights PSS-GOSO’s effectiveness in generating high-accuracy point clouds for complex mapping scenarios involving diverse data sources and sessions.
@@ 6.5 Evaluation on Proprietary Cross-platform Indoor-to-outdoor Dataset
Table 4: GCP Errors of different methods in the multiple platforms dataset [m]. The best results are in BOLD.
Sequence Initial (SLAM + Regis) (Xu et al., 2022) PSS-GOSO Pose-graph  (Koide et al., 2019) HBA (Liu et al., 2023)
@@
6.6 Time efficiency of PSS-GOSO
To evaluate the efficiency of PSS-GOSO and the impact of graph sparsification, we conducts an ablation study using a proprietary cross-platform dataset that includes both indoor and outdoor environments, which are listed in Table 5. Initially, we disabled the optimality-aware graph sparsification module within the GOSO framework, using the complete graph with all relationships included. This resulted in a raw graph with 65,926 edges and a processing time exceeding 4,800 seconds, highlighting the substantial computational burden. In contrast, when we enable the graph sparsification module in PSS-GOSO, the number of edges is reduced to 13,185 based on optimality criteria, significantly decreasing the processing time to 1,695 seconds. Despite the reduction in computational complexity, the accuracy of the final results remains high at 0.98 meters, comparable to the accuracy obtained with the fully connected graph. We also compare the processing time with the pose-graph-based method (Koide et al., 2019), which takes 1,218 seconds to get the final results. This comparison demonstrates that PSS-GOSO, through its graph sparsification module, effectively enhances processing efficiency while maintaining high accuracy, thus validating its effectiveness for large-scale LiDAR bundle adjustments and high-performance 3D mapping applications.
Table 5: Time efficiency analysis of the PSS-GOSO on the Cross-platform indoor-to-outdoor dataset.
@@ 6.6 Time efficiency of PSS-GOSO
7 Conclusion
In this paper, we propose PSS-GOSO, a robust, efficient, and scalable LiDAR Bundle Adjustment (LBA) method designed to enhance the accuracy of large-scale point clouds. To improve robustness, PSS-GOSO utilizes progressive spatial smoothing to extract reliable LiDAR correspondences for optimization. For increased efficiency, PSS-GOSO analyzes the graph structure of the raw bundle adjustment problem and selects the most critical edges to conserve computational resources. To address large-scale bundle adjustments, PSS-GOSO employs graph marginalization to decompose the problem into manageable subproblems. The performance of PSS-GOSO was evaluated across various platforms (handheld, legged, UGV, UAV) and diverse scenes, demonstrating its potential for high-accuracy 3D mapping. In the near future, we plan to integrate image data into the adjustment system to further enhance accuracy.
8 Acknowledgment
This research was conducted under project WP5 within the Delta-NTU Corporate Lab with funding support from A*STAR under its IAF-ICP programme (Grant no: I2201E0013) and Delta Electronics Inc.
9 Declaration of Interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix A Auxiliary optimization
To minimize Eq. (14), we use the auxiliary function $\bm{\Xi}$ introduced by Xu et al. (2011):
@@ Appendix A Auxiliary optimization
Appendix B Relative Pose Residual and Covariance from Pair-wise Point Cloud Registration
@@
Appendix C Jacobian of Progressive Spatial Smoothing (PSS) Factor

### v2

@@ VI-C Evaluation on Air-based Benchmark Dataset
Figure 10: The point clouds and dense graphs in MARS-LVIG [41] dataset. Edges of the graph are marked as yellow lines. The nodes of the graph are marked as green axes.
Figure 11: Absolute Position Error (APE) distribution of PSS-GOSO in MARS-LVIG [41] dataset.
VI-D Evaluation on Proprietary Large-scale Port Dataset
The 3D prior map is essential for automatic robot delivery in large-scale port environments, where the complexity of the scene and the high volume of data present significant challenges. We conduct a comprehensive evaluation using data collected from an Ouster OS1-64 laser scanner across a $400\text{m}\times 400\text{m}$ port area, supplemented by RTK-GNSS data for ground truth. Initial pose estimates are obtained using Fast-LIO and manual loop closure. The dense graph and point clouds generated by PSS-GOSO are illustrated in Fig. 12, and the Absolute Pose Error (APE) distribution is shown in Fig. 13. The experimental results, summarized in Table III, reveal that PSS-GOSO achieves an APE of 1.31 meters for the port UGV sequence. This performance significantly surpasses the initial APE of 8.63 meters, the APE of 4.33 meters from the pose-graph method [38], and the 3.31 meters APE from the hierarchical bundle adjustment (HBA) method [2]. These results highlight PSS-GOSO’s superior accuracy and effectiveness in handling large-scale, complex environments, demonstrating its capability to improve 3D mapping and pose estimation for automated robot delivery in ports.
TABLE III: Absolute Position Error (APE) of different methods in the large-scale port dataset [m]. The best results are in BOLD.
Sequence Initial (SLAM + Loop)[37] PSS-GOSO Pose-graph  [38] HBA [2]
Port UGV 8.63 1.31 4.33 3.31
@@ VI-D Evaluation on Proprietary Large-scale Port Dataset
VI-E Evaluation on Proprietary Cross-platform Indoor-to-outdoor Dataset
The 3D prior map is essential for applications such as outdoor-to-indoor surveillance and elderly health care, where accurate spatial information is crucial. Large-scale mapping tasks often involve cross-platform and multi-session data fusion, posing significant challenges for existing LBA methods. Traditional sequential splitting strategies, such as those used in HBA [2], are insufficient for addressing the complexities of multi-session and cross-platform data integration. In contrast, PSS-GOSO effectively tackles these challenges by partitioning the graph based on spatial relationships, allowing it to manage cross-platform and multi-session fusion tasks more efficiently. For our evaluation, initial pose estimates are derived using Fast-LIO and manual registration of submaps from different sessions. The dense graph and point clouds generated by PSS-GOSO are shown in Fig. 14. To assess the accuracy of the point clouds produced by various methods, ground control points (GCPs) are collected using Leica RTC360 survey scanners as shown in Fig. 15. The average errors of the GCPs, detailed in Table IV, indicate that PSS-GOSO reduced initial errors from 3.02 meters to 0.98 meters, demonstrating superior accuracy compared to existing pose-graph-based methods. This highlights PSS-GOSO’s effectiveness in generating high-accuracy point clouds for complex mapping scenarios involving diverse data sources and sessions. Moreover, the large-scale indoor-to-outdoor is used for the robot navigation as shown in Fig. 16. The real-time scanning data is matched with the prior map to get the accurate position for the robot, so the robot can navigate in large-scale complex environments for delivery and surveillance tasks. The demonstration video can be found on the project page.
@@ VI-E Evaluation on Proprietary Cross-platform Indoor-to-outdoor Dataset
Figure 16: Navigation of the robot in the large-scale indoor-to-outdoor dataset. Demonstration video can be found on the project page.
TABLE IV: GCP Errors of different methods in the multiple platforms dataset [m]. The best results are in BOLD.
Sequence Initial (SLAM + Regis) [37] PSS-GOSO Pose-graph  [38] HBA [2]
@@
VI-F Time efficiency of PSS-GOSO
To evaluate the efficiency of PSS-GOSO and the impact of graph sparsification, we conduct an ablation study using a proprietary cross-platform dataset that includes both indoor and outdoor environments, which are listed in Table V. Initially, we disabled the optimality-aware graph sparsification module within the GOSO framework, using the complete graph with all relationships included. This resulted in a raw graph with 65,926 edges and a processing time exceeding 4,800 seconds, highlighting the substantial computational burden. In contrast, when we enable the graph sparsification module in PSS-GOSO, the number of edges is reduced to 13,185 based on optimality criteria, significantly decreasing the processing time to 1,695 seconds. Despite the reduction in computational complexity, the accuracy of the final results remains high at 0.98 meters, comparable to the accuracy obtained with the fully connected graph. We also compare the processing time with the pose-graph-based method [38], which takes 1,218 seconds to get the final results. This comparison demonstrates that PSS-GOSO, through its graph sparsification module, effectively enhances processing efficiency while maintaining high accuracy, thus validating its effectiveness for large-scale LiDAR bundle adjustments and high-performance 3D mapping applications.
TABLE V: Time efficiency analysis of the PSS-GOSO on the Cross-platform indoor-to-outdoor dataset.
@@ VI-F Time efficiency of PSS-GOSO
VII Conclusion
In this paper, we propose PSS-GOSO, a robust, efficient, and scalable LiDAR Bundle Adjustment (LBA) method designed to enhance the accuracy of large-scale point clouds. To improve robustness, PSS-GOSO utilizes progressive spatial smoothing to extract reliable LiDAR correspondences for optimization. For increased efficiency, PSS-GOSO analyzes the graph structure of the raw bundle adjustment problem and selects the most critical edges to conserve computational resources. To address large-scale bundle adjustments, PSS-GOSO employs graph marginalization to decompose the problem into manageable subproblems. The performance of PSS-GOSO was evaluated across various platforms (handheld, legged, UGV, UAV) and diverse scenes, demonstrating its potential for high-accuracy 3D mapping. Optimized point clouds are successfully used as the prior map for robot last-mile delivery. In the near future, we plan to integrate the image data into the adjustment system to further enhance accuracy.
Appendix A: Auxiliary optimization
To minimize Eq. (14), we use the auxiliary function $\bm{\Xi}$ introduced by [28]:
@@ VII Conclusion
Appendix B: Relative Pose Residual and Covariance from Pair-wise Point Cloud Registration
@@
Appendix C: Jacobian of Progressive Spatial Smoothing (PSS) Factor